OpenStudy (anonymous):

Find the 6th term of the sequence in which t1=-5 and tn=-5tn-1

OpenStudy (kirbykirby):

Plug in n=6 :) \(\large t_6=-5t_{6-1}\) \(\large t_6 = -5t_5\) We don't know \(t_5\) though.. so continue the same process \(\large t_5=-5t_{5-1}\) \(\large t_5 = -5t_4\) Hopefully you see the pattern. It will decrease like this until you reach \(t_1\) since this is your initial value, and \(t_1=-5\) \(\large t_2=-5t_{2-1}\) \(\large t_2 = -5t_1=-5(-5)=25\) Then: \(\large t_3=-5t_{2}\) \(\large t_3 = -5(25)=-125\) etc.

OpenStudy (kirbykirby):

and climb back up to \(t_6\) doing that procedure